OpenStudy (evansanchez):

All of the following expressions are equivalent except _____. -5(x - 1) -5x + 5 5 - 5x (5 - 5)x

Directrix (directrix):

@evansanchez What does this equal when simplified: (5 - 5)x

OpenStudy (evansanchez):

0

OpenStudy (afro):

do distributive property

Directrix (directrix):

@evansanchez What does this equal when simplified: (5 - 5)x You are correct. (5-5)*x = 0*x = 0 The other options are all equal: -5(x - 1) -5x + 5 5 - 5x

Directrix (directrix):

This is the one that is not equivalent to the others because it collapses to zero. (5 - 5)x